# Sparsity Forcing: Reinforcing Token Sparsity of MLLMs

### v3

@@ Abstract
Sparse attention mechanisms aim to reduce computational overhead with minimal accuracy loss by selectively processing salient tokens. Despite their effectiveness, most methods merely exploit a model’s inherent sparsity and thus plateau at moderate budgets (about 50% token reduction), with little headroom to push budget lower without hurting accuracy. Other approaches attempt to enforce sparsity through trainable sparse attention or sharpness-inducing regularizers, but these either fix rigid patterns that ignore input and layer dynamics, or optimize proxy objectives without direct control over token budgets. In this paper, we explicitly reinforce token sparsity in well-posed multimodal large language models (MLLMs) through a simple RL-based post-training framework named Sparsity Forcing. Our method explores the efficiency-accuracy trade-off by running multiple rollouts with different token budgets, where both efficiency (token reduction ratio) and performance (answer correctness) are formulated as joint rewards. By contrasting rollouts within each group, the more efficient and correct answer is rewarded while less efficient or incorrect ones are penalized, thereby turning token saving into an end-to-end, inference-consistent optimization objective. Across thirteen image and video benchmarks, Sparsity Forcing raises token reduction ratio on Qwen2-VL/Qwen2.5-VL from 20% to 75% with minimal accuracy decline, significantly reducing long-context inference memory by up to 3 while speeding up decoding by up to 3.3.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (Chen et al., 2024c; Li et al., 2024a; Yang et al., 2024a) have achieved impressive results in tasks such as image captioning and visual question answering (Fu et al., 2023; Li et al., 2024b). However, when processing high-resolution images or long videos, the visual encoder often produces an excessive number of visual tokens, severely constraining the generative efficiency of MLLMs (Chen et al., 2024a). Existing sparse attention methods (Chen et al., 2024a; He et al., 2024; Jiang et al., 2024) exploit the sparsity of attention maps to prune redundant tokens, thereby reducing memory and latency. The core idea is that many tokens receive negligible attention and can be safely discarded, as in FastV (Chen et al., 2024a), which drops half of the visual tokens after the second layer with little accuracy loss. However, further budget reduction—for instance, to 20% or even 10% of the total tokens—remains challenging, since these approaches merely leverage naturally emergent sparsity of MLLMs rather than actively enforcing it, which fundamentally affects the achievable inference-time acceleration.
@@
In this paper, we propose a reinforcement learning-based post-training method, Sparsity Forcing , that directly optimizes the performance-efficiency trade-off of MLLMs via the Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For each vision-language query, we optimize a joint reward that increases with final-answer correctness and with token reduction ratio, turning token saving into an end-to-end objective rather than a proxy. Specifically, as shown in Figure 1, we treat an MLLM with sparse attention (e.g., Qwen2-VL+ZipVL (Wang et al., 2024; He et al., 2024)) as the policy model and the original MLLM with standard causal attention as the reference model; this anchoring stabilizes learning and preserves task fidelity, limiting accuracy loss even at high sparsity. To discover the smallest budget that preserves correctness, we execute multiple rollouts under adaptive token budgets to generate answers—a progressive budget sweep that dynamically tests whether low-salience tokens are actually needed for answer correctness. For each rollout, we compute the joint efficiency-performance reward, then contrast rewards within the group to define the advantages: rollouts that are both correct and more efficient tend to receive positive advantages, while less efficient or incorrect rollouts receive lower or negative advantages; then the policy is updated accordingly. Over training, this update progressively concentrates computation on the most informative tokens. Compared with previous SFT-based methods, our RL-based method applies an inference-aligned token pruning policy and KV-cache management for deployment, thereby yielding real end-to-end savings on MLLMs with minimal performance decline.
@@ 2 Related Work
Sparse attention of MLLMs. Sparse attention is designed to alleviate the computational constraints during deployment, which arise from the extended visual sequence. A line of work achieves it through token selection. ZipVL (He et al., 2024) proposed a dynamic ratio allocation strategy of important tokens, which is adaptively determined based on the layer-specific distribution of attention scores. VisionZip (Yang et al., 2024b) is a text-agnostic visual token compressor that selects dominant tokens by attention of the visual encoder and merges the rest by semantic similarity. SparseVLM (Zhang et al., 2024b) selects relevant text “raters” via attention, adaptively prunes visual tokens with a rank-based per-layer ratio, and recycles pruned tokens. SeerAttention (Gao et al., 2024) developed a self-distillation training scheme to efficiently train the AttnGate as a binary classification of token selection. Another line assigns sparsity patterns to attention heads. For example, Minference (Jiang et al., 2024) and MMinference (Li et al., 2025) determined the optimal pattern for each attention head and dynamically built sparse indices based on the assigned pattern during inference. However, these methods primarily leverage inherent attention sparsity at inference rather than actively encouraging token sparsity during optimization, which can lead to accuracy degradation under extremely low budgets. In this paper, we propose Sparsity Forcing, a post-training RL-based framework that explicitly encourages token sparsity by formulating token selection and answer correctness as an efficiency-performance reward-driven decision process that aligns with the actual inference pipeline.
@@ 3.2 Sparsity Forcing
Overview. As shown in Figure 1, we use an MLLM (e.g., Qwen2-VL (Wang et al., 2024)) with sparse attention (ZipVL (He et al., 2024)) as our policy model , and the same model with frozen parameters and standard causal attention as the reference model . For each question , we perform independent rollouts through the policy model to generate a set of distinct responses , where each response is generated with a random threshold , for . In addition, we record the token ratios associated with these responses. Then we compute the reward of each answer with a specially designed reward function that simultaneously considers the performance-efficiency trade-off. As illustrated in Figure 2, sweeping constitutes a progressive test of whether low-salience tokens are necessary: correctness changes across reveal the smallest budget that preserves accuracy. Such a more efficient and correct rollout receive a positive advantage in GRPO, while all other rollouts—either incorrect or less efficient—receive less or even a negative advantage. Finally, we use the corresponding advantages to update the policy model. Here are three key benefits of Sparsity Forcing: (1) We make the efficiency-performance trade-off an end-to-end objective rather than a proxy—GRPO jointly rewards token reduction and answer accuracy. (2) Multi-budget rollouts obtained by sweeping thresholds enable dynamic exploration of the minimum tokens needed for correctness across layers, inputs, and training dynamics, avoiding rigid, hand-crafted positive/negative labels for RL. (3) The training loop mirrors inference: the same token pruning policy and KV-cache management are applied during training and deployment, yielding predictable, budget-aware behavior across tasks and hardware.
@@
Inference is then carried out using only the important tokens in for each layer, which participate in subsequent self-attention during both prefilling and decoding, while unimportant tokens are skipped.
@@ 4 Experimental Setup
Implementation. We adopt QwenVL-series models (Qwen2-VL-7B (Wang et al., 2024), Qwen2.5-VL-3B (Yang et al., 2024a), and Qwen2.5-VL-7B) and LLaVA-series models (LLaVA-Video-7B (Zhang et al., 2024c)) as base MLLMs, owing to their strong adaptation to recent R1-based projects. For QwenVL-series calibration, we employ the Video-R1-260k (Feng et al., 2025) dataset, which covers both image and video modalities to ensure broad domain generalization. For LLaVA-Video, we use a 10k subset of LLaVA-Video-178k (Zhang et al., 2024c) to maintain its capability in the video domain. The parameter is set within the range of [0.94, 0.975] with a step size of 0.005 for training. We train our model on 8 NVIDIA A100 GPUs, limiting the number of video frames to 8 for efficiency. The Adam optimizer is used with a learning rate of 1e-6. In the KL divergence term of the GRPO algorithm, the hyperparameter is set to 0.04. During inference, Qwen2.5-VL and Qwen2-VL are evaluated on video benchmarks with approximately 120 frames corresponding to 16,384 tokens, while LLaVA-Video is evaluated with 110 frames corresponding to 20,240 tokens. Notably, during training we employ multiple-budget rollouts to explicitly probe the efficiency-performance trade-off, so that the model learns to enhance token sparsity. At inference, however, we simply fix to 0.975, which is the upper bound of the optimization range, to guarantee accuracy while still reaping the efficiency gains learned during training.
@@ 5 Main Results
| Full | 100 | 2303 | 62.2 | 84.0 | 64.5 | 70.2 |
| MOBA | 25.0 | 1906 | 58.6 | 77.3 | 62.6 | 66.2 |
| ZipVL | 61.7 | 2264 | 62.0 | 78.9 | 64.2 | 68.4 |
| Sharpness loss | 25.0 | 1965 | 59.6 | 77.0 | 63.7 | 66.8 |
| Sparsity Forcing | 26.4 | 2286 | 62.5 | 83.1 | 64.0 | 70.0 |
Performance comparison with post-training baseline methods in enhancing token sparsity. As shown in Table 3, Sparsity Forcing achieves a superior efficiency-performance trade-off compared to baseline methods on Qwen2.5-VL-7B. The standard-attention model reaches an average score of 70.0, while MOBA and sharpness loss, both operating at a 25% token ratio, suffer from noticeable drops to 66.2 and 66.8, respectively. ZipVL provides a moderate balance, obtaining 68.4 at a higher token ratio of 61.7%. In contrast, Sparsity Forcing attains 70.3 even with only 26.4% tokens, outperforming all baseline approaches on tasks such as MMStar (Chen et al., 2024b) and ChartQA (Masry et al., 2022). These results highlight the effectiveness of our RL-based optimization that reformulates the efficiency-performance tradeoff as explicit rewards.
Adjustment under different budgets. We further evaluate Sparsity Forcing under varying token budgets to illustrate its adaptability. As shown in Figure 3, Sparsity Forcing enables flexible budget adjustment across tasks and models. On commonsense benchmarks (Liu et al., 2024a), it maintains strong performance even under an extremely low budget of about 10% tokens. On OCR-related tasks (Liu et al., 2024c; Masry et al., 2022), it can further compress the token ratio while still preserving better accuracy than other methods, demonstrating its ability to focus on fine-grained textual regions. Moreover, this property generalizes well across both image and video benchmarks and different model backbones, consistently outperforming other sparse attention baselines under the same token ratio.
@@ 7 Conclusion
Limitations and future work. Our current multiple-budget rollout scheme executes answers sequentially for each because every budget induces a different retention mask and KV-cache trajectory, which precludes straightforward batching and slows training. In our future work, we will address this limitation to enable easier and more efficient training. Moreover, beyond token budgets, we also plan to extend reinforced efficiency to broader applications and objectives—such as hardware-aware targets (latency/memory/energy), tool-/retrieval-call budgets, head/layer/MoE expert gating, and KV/activation quantization and retention.

### v4

@@ Abstract
Sparse attention mechanisms aim to reduce computational overhead with minimal accuracy loss by selectively processing salient tokens. Despite their effectiveness, most methods merely exploit a model’s inherent sparsity and thus plateau at moderate budgets (about 50% token reduction), with little headroom to push budget lower without hurting accuracy. Other approaches attempt to enforce sparsity through trainable sparse attention or sharpness-inducing regularizers, but these either fix rigid patterns that ignore input and layer dynamics, or optimize proxy objectives without direct control over token budgets. In this paper, we explicitly reinforce token sparsity in well-posed multimodal large language models (MLLMs) through a simple RL-based post-training framework named Sparsity Forcing. Our method explores the efficiency-accuracy trade-off by running multiple rollouts with different token budgets, where both efficiency (token reduction ratio) and performance (answer correctness) are formulated as joint rewards. By contrasting rollouts within each group, the more efficient and correct answers are rewarded while less efficient or incorrect ones are penalized, thereby turning token saving into an end-to-end, inference-consistent optimization objective. Across thirteen image and video benchmarks, Sparsity Forcing raises token reduction ratio on Qwen2-VL/Qwen2.5-VL from 20% to 75% with minimal accuracy decline, significantly reducing long-context inference memory by up to 3 while speeding up decoding by up to 3.3.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (Chen et al., 2024c; Li et al., 2024a; Yang et al., 2024a) have achieved impressive results in tasks such as image captioning and visual question answering (Fu et al., 2023; Li et al., 2024b). However, when processing high-resolution images or long videos, the visual encoder often produces an excessive number of visual tokens, severely constraining the generative efficiency of MLLMs (Chen et al., 2024a). Existing sparse attention methods (Chen et al., 2024a; He et al., 2024a; Jiang et al., 2024) exploit the sparsity of attention maps to prune redundant tokens, thereby reducing memory and latency. The core idea is that many tokens receive negligible attention and can be safely discarded, as in FastV (Chen et al., 2024a), which drops half of the visual tokens after the second layer with little accuracy loss. However, further budget reduction—for instance, to 20% or even 10% of the total tokens—remains challenging, since these approaches merely leverage naturally emergent sparsity of MLLMs rather than actively enforcing it, which fundamentally affects the achievable inference-time acceleration.
@@
In this paper, we propose a reinforcement learning-based post-training method, Sparsity Forcing , that directly optimizes the performance-efficiency trade-off of MLLMs via the Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For each vision-language query, we optimize a joint reward that increases with final-answer correctness and with token reduction ratio, turning token saving into an end-to-end objective rather than a proxy. Specifically, as shown in Figure 1, we treat an MLLM with sparse attention (e.g., Qwen2-VL+ZipVL (Wang et al., 2024; He et al., 2024a)) as the policy model and the original MLLM with standard causal attention as the reference model; this anchoring stabilizes learning and preserves task fidelity, limiting accuracy loss even at high sparsity. To discover the smallest budget that preserves correctness, we execute multiple rollouts under adaptive token budgets to generate answers—a progressive budget sweep that dynamically tests whether low-salience tokens are actually needed for answer correctness. For each rollout, we compute the joint efficiency-performance reward, then contrast rewards within the group to define the advantages: rollouts that are both correct and more efficient tend to receive positive advantages, while less efficient or incorrect rollouts receive lower or negative advantages; then the policy is updated accordingly. Over training, this update progressively concentrates computation on the most informative tokens. Compared with previous SFT-based methods, our RL-based method applies an inference-aligned token pruning policy and KV-cache management for deployment, thereby yielding real end-to-end savings on MLLMs with minimal performance decline.
@@ 2 Related Work
Sparse attention of MLLMs. Sparse attention is designed to alleviate the computational constraints during deployment, which arise from the extended visual sequence. A line of work achieves it through token selection. ZipVL (He et al., 2024a) proposed a dynamic ratio allocation strategy of important tokens, which is adaptively determined based on the layer-specific distribution of attention scores. VisionZip (Yang et al., 2024b) is a text-agnostic visual token compressor that selects dominant tokens by attention of the visual encoder and merges the rest by semantic similarity. SparseVLM (Zhang et al., 2024b) selects relevant text “raters” via attention, adaptively prunes visual tokens with a rank-based per-layer ratio, and recycles pruned tokens. SeerAttention (Gao et al., 2024) developed a self-distillation training scheme to efficiently train the AttnGate as a binary classification of token selection. Another line assigns sparsity patterns to attention heads. For example, Minference (Jiang et al., 2024) and MMinference (Li et al., 2025) determined the optimal pattern for each attention head and dynamically built sparse indices based on the assigned pattern during inference. However, these methods primarily leverage inherent attention sparsity at inference rather than actively encouraging token sparsity during optimization, which can lead to accuracy degradation under extremely low budgets. In this paper, we propose Sparsity Forcing, a post-training RL-based framework that explicitly encourages token sparsity by formulating token selection and answer correctness as an efficiency-performance reward-driven decision process that aligns with the actual inference pipeline.
@@ 3.2 Sparsity Forcing
Overview. As shown in Figure 1, we use an MLLM (e.g., Qwen2-VL (Wang et al., 2024)) with sparse attention (ZipVL (He et al., 2024a)) as our policy model , and the same model with frozen parameters and standard causal attention as the reference model . For each question , we perform independent rollouts through the policy model to generate a set of distinct responses , where each response is generated with a random threshold , for . In addition, we record the token ratios associated with these responses. Then we compute the reward of each answer with a specially designed reward function that simultaneously considers the performance-efficiency trade-off. As illustrated in Figure 2, sweeping constitutes a progressive test of whether low-salience tokens are necessary: correctness changes across reveal the smallest budget that preserves accuracy. Such a more efficient and correct rollout receive a positive advantage in GRPO, while all other rollouts—either incorrect or less efficient—receive less or even a negative advantage. Finally, we use the corresponding advantages to update the policy model. Here are three key benefits of Sparsity Forcing: (1) We make the efficiency-performance trade-off an end-to-end objective rather than a proxy—GRPO jointly rewards token reduction and answer accuracy. (2) Multi-budget rollouts obtained by sweeping thresholds enable dynamic exploration of the minimum tokens needed for correctness across layers, inputs, and training dynamics, avoiding rigid, hand-crafted positive/negative labels for RL. (3) The training loop mirrors inference: the same token pruning policy and KV-cache management are applied during training and deployment, yielding predictable, budget-aware behavior across tasks and hardware.
@@
Inference is then carried out using only the important tokens in for each layer, which participate in subsequent self-attention during both prefilling and decoding, while unimportant tokens are skipped. For implementation, we keep a full KV cache and perform dynamic sparse token selection only at decoding time by fetching the corresponding KV entries on the fly (Kwon et al., 2023).
@@ 4 Experimental Setup
Implementation. We adopt QwenVL-series models (Qwen2-VL-7B (Wang et al., 2024), Qwen2.5-VL-3B (Yang et al., 2024a), and Qwen2.5-VL-7B) and LLaVA-series models (LLaVA-Video-7B (Zhang et al., 2024c)) as base MLLMs, owing to their strong adaptation to recent R1-based projects. For QwenVL-series calibration, we employ the Video-R1-260k (Feng et al., 2025) dataset, which covers both image and video modalities to ensure broad domain generalization. For LLaVA-Video, we use a 10k subset of LLaVA-Video-178k (Zhang et al., 2024c) to maintain its capability in the video domain. The parameter is set within the range of [0.94, 0.975] with a step size of 0.005 for training. We train our model on 8 NVIDIA A100 GPUs with 883 GPU hours for Qwen2.5VL-7b and 164 GPU hours for LLaVA-Video-7b, limiting the number of video frames to 8 for efficiency. The Adam optimizer is used with a learning rate of 1e-6. In the KL divergence term of the GRPO algorithm, the hyperparameter is set to 0.04. During inference, Qwen2.5-VL and Qwen2-VL are evaluated on video benchmarks with approximately 120 frames corresponding to 16,384 tokens, while LLaVA-Video is evaluated with 110 frames corresponding to 20,240 tokens. Notably, during training we employ multiple-budget rollouts to explicitly probe the efficiency-performance trade-off, so that the model learns to enhance token sparsity. At inference, however, we simply fix to 0.975, which is the upper bound of the optimization range, to guarantee accuracy while still reaping the efficiency gains learned during training.
@@ 5 Main Results
| Full | 100 | 2303 | 62.2 | 84.0 | 64.5 | 73.2 |
| MOBA | 25.0 | 1906 | 58.6 | 77.3 | 62.6 | 66.6 |
| ZipVL | 61.7 | 2264 | 62.0 | 78.9 | 64.2 | 71.5 |
| Sharpness loss | 25.0 | 1965 | 59.6 | 77.0 | 63.7 | 67.6 |
| Sparsity Forcing | 26.4 | 2286 | 62.5 | 83.1 | 64.0 | 72.8 |
Performance comparison with post-training baseline methods in enhancing token sparsity. As shown in Table 3, Sparsity Forcing achieves a superior efficiency-performance trade-off compared to baseline methods on Qwen2.5-VL-7B. The standard-attention model reaches an average score of 73.2, while MOBA and sharpness loss, both operating at a 25% token ratio, suffer from noticeable drops to 66.6 and 67.8, respectively. ZipVL provides a moderate balance, obtaining 71.5 at a higher token ratio of 61.7%. In contrast, Sparsity Forcing attains 72.8 even with only 26.4% tokens, outperforming all baseline approaches on tasks such as MMStar (Chen et al., 2024b) and ChartQA (Masry et al., 2022). These results highlight the effectiveness of our RL-based optimization that reformulates the efficiency-performance tradeoff as explicit rewards.
Adjustment under different budgets. We further evaluate Sparsity Forcing under varying token budgets to illustrate its adaptability. As shown in Figure 3, Sparsity Forcing enables flexible budget adjustment across tasks and models. On commonsense benchmarks (Liu et al., 2024a), it maintains strong performance even under an extremely low budget of about 10% tokens. On OCR-related tasks (Liu et al., 2024c; Masry et al., 2022), it can further compress token ratio while still preserving better accuracy than other methods, demonstrating its ability to focus on fine-grained textual regions. Moreover, this property generalizes well across both image and video benchmarks and different model backbones, outperforming other sparse attention baselines under the same token ratio.
@@ 7 Conclusion
Limitations and future work. Our current work mainly focuses on single-turn hardware-agnostic token sparsity. In our future work, we plan to extend reinforced efficiency to broader applications and objectives—such as hardware-aware targets (latency/memory/energy), multi-turn dialogue, tool-/retrieval-call budgets, head/layer/MoE expert gating, and KV/activation quantization and retention.
Appendix A Appendix
A.1 Algorithm of our dynamic sparse attention
As shown in Algorithms 1 and 2, we summarize the sparse attention mechanisms used in our method for the prefill and decoding.
A.2 Implementation of Attention Sharpness Loss
Following prior works such as MOBA (Lu et al., 2025) and NSA (Yuan et al., 2025), we adopt a block-wise probing strategy by applying average pooling over queries and keys with block size to approximate the attention map. Given the visual input with query and key , we first compress the sequence via average pooling:
where is applied along the sequence dimension, yielding compressed representations . The block-wise attention map is then computed as
where denotes the function and is the causal attention mask.
To promote sparsity, we encourage each query to concentrate on a few salient keys by introducing an sharpness-inducing regularization:
where denotes the -th row of . This loss increases the peak mass of each query’s attention distribution, enforcing sharper focus and reducing redundancy among tokens. At inference time, we apply sparse attention by retaining only the top 25% salient tokens according to .
A.3 Implementation of Sparsity Forcing with different sparse attention
To implement Sparsity Forcing with different sparse attention mechanisms, we first compute the accumulated attention score of each token as in Eq. 4, and normalize it to using Eq. 6, since early tokens accumulate higher values than later ones (He et al., 2024b). For top- sparse attention, we retain the most salient tokens ranked by for subsequent attention. For threshold-based sparse attention with threshold , we keep all tokens with .
During training, we adopt multiple budget rollouts to probe the efficiency-accuracy trade-off. For top- sampling, we set within [25%, 50%] at intervals of 5%, where 25% reflects the target efficiency budget while 50% corresponds to the accuracy-preserving range of most sparse attention methods. For threshold-based attention, we empirically vary within [, ] with a step of 0.002. At inference time, we fix the budget by retaining the top 25% salient tokens for top- sparse attention and set for threshold-based sparse attention.
A.4 Failure case
These failure cases in Figure 7 highlight a key limitation of token pruning, especially with enhanced sparsity: while it effectively reduces redundancy, it may inadvertently disrupt cross-frame correspondences that are critical for spatial reasoning. Unlike commonsense reasoning, which often relies on text-grounded cues, spatial understanding demands alignment of visual features across frames—such as consistent tracking of object positions, relative distances, and scene layouts.
A.5 More Results and Discussion
| Efficiency reward | Ratio | VideoMME | MLVU | VideoMMMU | PerceptionTest | EgoSchema | TempCompass | Avg. |
| Token reduction ratio | 29.6 | 64.1/70.2 | 70.3 | 35.2 | 66.0 | 57.0 | 66.2 | 61.3 |
| Latency + token reduction ratio | 30.5 | 63.9/70.3 | 70.4 | 35.0 | 66.4 | 56.4 | 66.0 | 61.2 |
Hardware efficiency reward. Hardware latency is an important practical metric. We therefore conducted an additional ablation in which we explicitly include measured hardware latency in the efficiency reward. Concretely, we compare two variants on LLaVA-Video-7B: (i) using only the token reduction ratio as the efficiency term, and (ii) using a combined efficiency reward that depends on both latency and token reduction ratio. As shown in Table 6, incorporating latency into the reward does not yield noticeable benefits: both variants achieve almost identical accuracy (61.3 vs. 61.2 Avg.) with very similar token ratios. These results suggest that the token reduction ratio is already a sufficiently good surrogate for hardware efficiency in our setting.
| Method | Wall-clock training time (hour) | Latency(s) |
| MOBA | 75.6 | 11.6 |
| ZipVL | 81.0 | 18.7 |
| Sharpness loss | 88.4 | 11.9 |
| Sparsity Forcing (ours) | 110.4 | 12.0 |
Training and rollout cost comparison with baseline methods. As shown in Table 7, we note our method is moderately slower than other SFT-based baselines but still within an acceptable range. This is mainly because GRPO-based fine-tuning requires the MLLM to generate multiple rollouts for each sample, whereas other methods only perform standard fine-tuning. In terms of rollout efficiency, our method is 1.6× faster than the SFT-based ZipVL and remains comparable to other methods, since MOBA and sharpness loss use manually fixed token budgets. However, as shown in Table 3, our method achieves a superior performance–efficiency tradeoff compared with SFT-based counterparts on both image and video benchmarks, making the additional training cost worthwhile.
| Method | Ratio | MME | MMStar | ChartQA | VideoMME | Avg |
| Full | 100 | 2303 | 62.2 | 84.0 | 64.5 | 70.2 |
| MOBA | 25.0 | 1906 | 58.6 | 77.3 | 62.6 | 66.2 |
| ZipVL | 61.7 | 2264 | 62.0 | 78.9 | 64.2 | 68.4 |
| Sharpness loss | 25.0 | 1965 | 59.6 | 77.0 | 63.7 | 66.8 |
| Temperature adjustment | 25.0 | 1960 | 59.8 | 78.6 | 61.3 | 67.4 |
| Sparsity Forcing(ours) | 26.4 | 2286 | 62.5 | 83.1 | 64.0 | 70.0 |
Additional comparison with temperature adjustment baseline method. Intuitively, temperature adjustment in the Softmax function can also sharpen the attention distribution. As shown in Table 8, temperature-based sharpening performs better than the sharpness-loss baseline, yet it still underperforms our method by 2.6%.